# Dynamic Sub-graph Distillation for Robust Semi-supervised Continual Learning

### v1

@@ Introduction
Figure 2: Baselines that combine CL and SSL methods to tackle SSCL. (a) iCaRL&Fix_DistillLabel: distillation on labeled data. iCaRL&Fix_DistillALL: distillation on the entire memory buffer. iCaRL&Fix_GTUnlabel: correct the pseudo-labels of unlabeled data in the memory buffer to ground truth. (b) DER&Fix_New(Old)_pseudo: apply the predictions of the current network as pseudo labels $\bm{p}^{\mathcal{A}}$. DER&Fix_Old_pseudo: use the previous prediction as $\bm{p}^{\mathcal{A}}$. (c-d) The distribution bias of fully and semi-supervised settings.
@@
Semi-supervised Continual Learning (SSCL) considers a more realistic continual task stream that only a limited number of samples are annotated. The success of SSL allows a general effort to exploit unlabeled data to improve performance throughout the entire stage. CNNL (Baucum et al. 2017) fine-tunes its incremental learner by generating the pseudo-labels of unlabeled to enable self-training. DistillMatch (Smith et al. 2021) employs knowledge distillation with prediction consistency on unlabeled data and optimizes an out-of-distribution detector to identify task-specific representations. Pseudo Gradient Learners (Luo et al. 2022) proposes a gradient learner from labeled data to predict gradients on unlabeled data to avoid the risk of pseudo labels.
@@ Methods
The research of SSCL amounts to learning an ordered set of $T$ tasks that exhibit different data distribution $\mathcal{D}^{t}$. The data of each task sampled i.i.d. from $\mathcal{D}^{t}$ with few annotations $D^{t}=\{(X^{t}_{l},Y^{t}_{l}),X^{t}_{u}\}$, where $X^{t}_{l}$ and $X^{t}_{u}$ represents the labeled and unlabeled data, respectively. In this paper, we consider class continual learning so that for any two tasks $Y^{s}\cap Y^{t}=\emptyset$.
Each task is a specific semi-supervised learning process, which attempts to find a model $f:X^{t}\rightarrow Y^{t}$ to map both labeled samples and unlabeled samples to the target space. After learning each task, the learned model should perform well on the previous tasks even without experience id by involving distillation loss or regularization constraints. In summary, the whole object can be formulated as follows:
@@ A Systematic Study of SSCL
[…]

### v2

@@ Introduction
Figure 2: Baselines that combine CL and SSL methods to tackle SSCL. (a) DistillLabel: distillation on labeled data. DistillALL: apply distillation loss on the entire memory buffer. GTUnlabel: correct the pseudo-labels of unlabeled data in the memory buffer to ground truth. (b) New(Old)_pseudo: apply the predictions of the current(previous) network as pseudo labels $\bm{p}^{\mathcal{A}}$. (c-d) The distribution bias of fully and semi-supervised settings.
@@
Semi-supervised Continual Learning (SSCL) considers a more realistic continual task stream that only a limited number of samples are annotated. The success of SSL allows a general effort to exploit unlabeled data to improve performance throughout the entire stage. CNNL (Baucum et al. 2017) fine-tunes its incremental learner by generating the pseudo-labels of unlabeled to enable self-training. DistillMatch (Smith et al. 2021) employs knowledge distillation with prediction consistency on unlabeled data and optimizes an out-of-distribution detector to identify task-specific representations. NNCSL (Kang et al. 2023) proposes class-instance similarity distillation to preserve relation information. Pseudo Gradient Learners (Luo et al. 2022) proposes a gradient learner from labeled data to predict gradients on unlabeled data to avoid the risk of pseudo labels.
@@ Methods
The research of SSCL amounts to learning an ordered set of $T$ tasks that exhibit different data distribution $\mathcal{D}^{t}$. The data of each task sampled i.i.d. from $\mathcal{D}^{t}$ with few annotations $D^{t}=\{(X^{t}_{l},Y^{t}_{l}),X^{t}_{u}\}$, where $X^{t}_{l}$ and $X^{t}_{u}$ represents the labeled and unlabeled data, respectively. In this paper, we consider class continual learning so that for any two tasks $Y^{s}\cap Y^{t}=\emptyset$. Each task is a specific semi-supervised learning process, which attempts to find a model $f:X^{t}\rightarrow Y^{t}$ to map both labeled samples and unlabeled samples to the target space. Furthermore, the learned model should perform well on the previous tasks even without experience id by involving distillation constraints. Therefore, the whole object can be formulated as follows:
@@ A Systematic Study of SSCL
Table 1: Average incremental accuracy and the last incremental accuracy. NNSCL5120 is the state-of-art method that conducts experiments on various label ratios, where the subscript 5120 represents the memory buffer size. The percentage values in brackets and subscripts are label ratios. Improvement represents the progress of our method DSGD compared to corresponding baselines. JointTrain means learning the entire dataset at once.